OpenStudy (anonymous):

(x-3)^2 +(y+4)^2=25 in polar form please help!!! Here are the choices a) r=6cos(theta)-8sin(theta) b) r=8sin(theta)-6cos(theta) c) r=7/(cos(theta)*sin(theta) d) r=4/(cos(theta)*sin(theta)

OpenStudy (anonymous):

answer is a. i got it by expanding the brackets, then plugging in these 3 equations: x^2 + y^2 = r^2 x = rcos(theta) y = rsin(theta)

OpenStudy (anonymous):

Thank you!!